OpenStudy (anonymous):

Value of \[\tan (\frac{ 1 }{ 2 } \cos ^{-1}\frac{ \sqrt{5} }{3 })\]

OpenStudy (ash2326):

@Yahoo! first find the value of \[\tan ( \cos ^{-1}\frac{ \sqrt{5} }{3 })\] could you find it?

OpenStudy (anonymous):

\[\tan ( \tan ^{-1} \frac{ 2 }{ \sqrt{5} })\]

OpenStudy (ash2326):

@Yahoo! I can see that you have closed your previous question before it was done. Please don't do this. It's against the code of conduct http://openstudy.com/code-of-conduct Thanks

OpenStudy (anonymous):

lol...@ash2326 i got...that answer....and was waiting...for...Sauravs method..

OpenStudy (anonymous):

@sauravshakya Any Idea

OpenStudy (anonymous):

@ash2326 ...()

OpenStudy (anonymous):

CANT WE USE CALCULATOR?

OpenStudy (anonymous):

No Calculator.)

OpenStudy (ash2326):

let \[x=\cos^{-1} (\frac{\sqrt 5}{3})\] so we got \[\tan x=\frac{2}{\sqrt 5}\] now find \[\tan (\frac x 2)\] we know \[\large \tan x=\frac{2\tan \frac x 2}{1-\tan^2 \frac x2}\] now find tan x/2 from this

OpenStudy (anonymous):

i think....this may be little easier: \[\cos ^{-1} \frac{ \sqrt{5} }{ 3 } = 2x\]

OpenStudy (ash2326):

it doesn't matter, you have to solve a quadratic at last. Substitute tan x/2 or tan x= y

OpenStudy (anonymous):

i got this...)....

OpenStudy (anonymous):

Which one shuld i choose

OpenStudy (anonymous):

Both of them r in my option

OpenStudy (ash2326):

the one which is positive

OpenStudy (anonymous):

Why...

OpenStudy (anonymous):

\[\frac{ \pm3 - \sqrt{5} }{ 2 }\]

OpenStudy (ash2326):

since \[x=\cos^{-1}\frac{\sqrt{5}}{3} \] \[\text{x will lie in the range 0 to}\frac{\pi}{2} \text{since }\frac{\sqrt{5}}{3} \text{is positive }\] so \[\text{x/2 will also lie in the range 0 to}\frac{\pi}{2}\] \[\large \text{tangent is positive in this range :)}\]

OpenStudy (anonymous):

Got.....it............thxxxxx

OpenStudy (ash2326):

welcome :)